OpenStudy (anonymous):

x^2 = -4x + 5, Quadratic problem, completing the square.

OpenStudy (anonymous):

x^2+4x - 5=0 add and subtract square of half of the coefficient of x i.e (4/2)^2 = 4 So x^2 + 4x +4 -4 -5 = 0 (x^2 + 4x +2^2) -4 -5 = 0 you know a^2 + 2ab + b^2 = (a+b)^2 a = x and b = 2 we have (x+2)^2 -9 = 0

OpenStudy (anonymous):

Ohhh okay, I was struggling to put the 4x - 5 to the other side, one question what is the i.e stand for?

OpenStudy (anonymous):

i.e stands for "that is"

OpenStudy (anonymous):

okay thank you :D